OpenStudy (anonymous):

find y for the linear equations 2x-3y=6 and find y for -10x=5y

OpenStudy (wannabegurl):

duno

OpenStudy (anonymous):

rearrange for y. 1) 2x-3y=6 2x -6 = 3y 2/3x-6/3 = y y=2/3x -2 2) -10x = 5y try this one yourself.

OpenStudy (anonymous):

ok wannabeagirl then go away

OpenStudy (wannabegurl):

\[gj\]

OpenStudy (wannabegurl):

retriceole

OpenStudy (anonymous):

@jayds i got for 1 x-1.3

OpenStudy (anonymous):

i really don t under stand

OpenStudy (anonymous):

find y means that you have to isolate y, do you see how I did it above? we want y = ............... Lets use an example say 5x + y = 9 I want to find "x" this time not y, so I just bring the y to the right hand side and bring the 9 to the left hand side. 5x - 9 = -y y = -5x + 9 do you understand?

OpenStudy (anonymous):

so if i put -1 in the place of x what will it be

OpenStudy (anonymous):

sorry, I do not understand what you are trying to say.

OpenStudy (anonymous):

so instead of y=2/3x -2 what will it equal if \[ y=2/3\times-1 -2\]

OpenStudy (anonymous):

you know what im saying

OpenStudy (anonymous):

ok, substitution of x=-1 inside the equation is what u mean. then you just solve it. y=2/3 x -1 -2 y= -2/3 -2 y=-3.667

OpenStudy (anonymous):

okkkkkkkkkkkk now i understand thanks so much!

OpenStudy (anonymous):

it's ok, medal thanks :P